# The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals

### v2

@@
Abstract.
The voluntary donation of private health information for altruistic purposes, such as supporting research advancements, is a common practice. However, concerns about data misuse and leakage may deter people from donating their information. Privacy Enhancement Technologies (PETs) aim to alleviate these concerns and in turn allow for safe and private data sharing. This study conducts a vignette survey ($N=494$) with participants recruited from Prolific to examine the willingness of US-based people to donate medical data for developing new treatments under four general guarantees offered across PETs: data expiration, anonymization, purpose restriction, and access control. The study explores two mechanisms for verifying these guarantees: self-auditing and expert auditing, and controls for the impact of confounds including demographics and two types of data collectors: for-profit and non-profit institutions.
@@ Abstract.
Privacy Guarantee, Health Data Donation, Privacy Expectation, Auditing
ccs: Security and privacy Social aspects of security and privacy
1. Introduction
@@
The altruistic use of personal health data holds immense potential for societal benefit. For example, the public health sector can leverage individuals’ medical histories and health data to analyze epidemiological trends, enhance disease surveillance and risk prediction, diagnose rare and emerging diseases, and accelerate the development of new treatments (Bietz et al., 2015; Aitken et al., 2016; Bartlett et al., 2018; Greene et al., 2019; Kim et al., 2015; Trinidad et al., 2020). Beyond these applications, there are significant economic advantages. The European Commission estimates that innovations enabling greater data sharing and availability could save billions of euros within the EU health sector (Commission, [n. d.]). As a result, the demand for personal health data continues to grow (Wilbanks and Friend, 2016; Bietz et al., 2015), particularly in light of the recent surge in large language models (LLMs), whose effectiveness heavily depends on the data they are trained on.
Upon the acceptance of data-sharing agreements, personal health data is routinely collected by healthcare providers (Bartlett et al., 2018; Kim et al., 2015) as well as personal mobile apps and devices (Greene et al., 2019). This data is often accessed by third parties through data-sharing agreements that govern data donation practices (Trinidad et al., 2020). However, people frequently self-report that privacy concerns deter them from donating their sensitive information (Redmiles, 2020; Voigt et al., 2020; Silber et al., 2022; Greene et al., 2019; Kim et al., 2015). For instance, many individuals express discomfort with hospital policies that permit data-sharing agreements with third parties without obtaining explicit consent from the data owners (Lewis, 2019).
In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage (Popa et al., 2011; Pappas et al., 2014), data processing (Arasu et al., 2013; Bajaj and Sion, 2011; Baumann et al., 2015; Gupta et al., 2019; Chowdhury et al., 2019; Al-Rubaie and Chang, 2019; Gaballah et al., 2024), and machine learning (ML) training (Wu et al., [n. d.]; De Viti et al., 2022). For example, CoVault (De Viti et al., 2022) enables multiparty analytics for epidemics-related datasets while preserving data confidentiality, and Anonify (Gaballah et al., 2024) supports decentralized, dual-level anonymity for secure medical data donation.
However, in order for such PETs to impact the availability of data, the guarantees they offer must impact user privacy concerns significantly enough to alter their willingness to donate their data.
@@
In order to evaluate whether the guarantees offered by PETs ultimately impact people’s willingness to donate their data, at least three steps are necessary. First, we must be able to effectively explain guarantees (either general guarantees or those of a particular PET) to the people they aim to protect: end-users. Prior work has focused on doing so for particular PETs, chiefly differential privacy (DP) and end-to-end encryption (E2EE) (Cummings et al., 2021; Abu-Salma et al., 2018; Akgul et al., 2021; Bai, 2019; De Luca et al., 2016; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023). Second, using these explanations, we must evaluate whether the guarantees impact people’s privacy concerns (or expectations), and in turn, their willingness to share data. Prior work has done such evaluations using self-report studies focused on individual PETs, again primarily DP (Cummings et al., 2021; Bullek et al., 2017; Nanayakkara et al., 2023) or E2EE (Dechand et al., 2019; Bai et al., 2020; Abu-Salma et al., 2018)). Third, while rare, we must validate those results in the field to confirm that the effects observed replicate when studying actual vs. intended behavior (e.g., (Dooley et al., 2022b; Acquisti et al., 2013)).
Instead of focusing on a singular PET and it’s accompanying guarantees, we focus on addressing steps one and two for the general security and privacy guarantees that recur across PETs:
  1. PG(1):
    Anonymization (Murthy et al., 2019; Shahid et al., 2022): data is not linkable to its owner (as defined in (Chevrier et al., 2019));
  2. PG(2):
    Access control (Greene et al., 2019; Hu et al., 2006; Ni et al., 2010; Tourani et al., 2017): data is accessible only by authorized people, specified in the data collection agreement;
  3. PG(3):
    Data expiration (Politou et al., 2018; Sarkar et al., 2018; Perlman, 2005): data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
  4. PG(4):
    Purpose restriction (De Viti et al., 2022; Byun et al., 2005): data is only used for the stated collection purpose, specified in the data collection agreement;
While previous research has explored how the combined presence of guarantees PG(1)-PG(4) affects data-sharing intentions (Aitken et al., 2016; Mori, 2016), our study takes a different approach by focusing on the impact of each guarantee individually. Our objective is to identify which specific protections or factors most significantly influence privacy expectations and, in turn, reported data-sharing intentions.
Additionally, we investigate the role of auditing or verification processes increasingly introduced to produce public auditing results for PETs, specifically:
  1. AG(1):
    Expert auditing (Li et al., 2015; Wang et al., 2011): engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
  2. AG(2):
    Self auditing (Holt and Seamons, 2005; Chen et al., 2022): granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
@@ 2. AG(2):
We examine the role of these PET guarantees and their audits in shaping privacy expectations and influencing data-sharing intentions within a specific context used in prior work (see e.g., (Cummings et al., 2021; Nanayakkara et al., 2023)): medical data donation. We ask:
@@ 1. RQ(1):
    How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?
@@ 2. RQ(2):
To address these questions, we conduct a vignette survey ($n=494$) following the best practice methodology established by prior work (Cummings et al., 2021; Abu-Salma et al., 2018; Xiong et al., 2020). Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with information on how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)). Furthermore, we isolate the effect of PG(1)–PG(4) and AG(1)–AG(2) from external confounding factors identified by prior work:
  • The data-collection entity (Trinidad et al., 2020; Hendricks-Sturrup and Lu, 2023): whether the collecting entity is a for-profit or a non-profit organization;
  • Socio-demographics (Hargittai et al., 2020; Dooley et al., 2022a; Trinidad et al., 2020): age, gender, education, donation history, and technical background.
@@
Furthermore, while the technical community has suggested external audits as a mechanism to increase trust in PET implementation, our initial inquiry suggests that more work is needed to align the perceptions of the experts and end users about the effectiveness of such audits. In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
We argue that it is critical for non-profit entities to rigorously implement data privacy measures, as any future data leak could lead to a significant loss of trust. Prior research (Acquisti et al., 2013) underscores that users place greater value on maintaining their expected privacy than on gaining additional privacy they did not initially anticipate. In contrast, we highlight the risk of for-profit entities engaging in ”privacy washing” (Cirucci, 2024), where statements about PETs are used to artificially raise privacy expectations to encourage data collection. At the same time, our findings reveal that respondents are perceptive of the limitations in general PET guarantees, particularly concerning protections against data breaches. This underscores the need for future research to explore how to effectively communicate the potential of stronger, emerging PET guarantees and address the skepticism of end-users.
2. Related Work
In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of privacy-enhancing technologies (PETs), and educational or explanatory strategies related to PETs.
@@ 2. Related Work
Data donation.Data donationData donation. The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments. Indeed, the COVID-19 pandemic has further intensified the demand for personal health data (Scott et al., 2020). Despite the importance of this data, its sensitive nature poses significant obstacles to its collection (Fahey and Hino, 2020). Prior work indicates that individuals are generally willing to donate their data for altruistic purposes (Pilgrim and Bohnet-Joschko, 2022; Aitken et al., 2016; Seberger and Patil, 2021), though there is a notable reluctance when it comes to their health data specifically (Middleton et al., 2020; Mori, 2016). However, willingness to donate personal health data increases if the donor or their close family members are directly affected by the disease under investigation (Goodman et al., 2017; Bartlett et al., 2018).
@@
Privacy concerns.Privacy concernsPrivacy concerns. Concerns about the privacy and the misuse of donated data are prominent among potential donors (Mori, 2016). These concerns include the risk of being identified, discriminated against, or having their personal sensitive data misused or leaked (Weitzman et al., 2012; Aitken et al., 2016; Kim et al., 2015; McCormick and Hopkins, 2021). A common source of these concerns is distrust of the receiving entity (Weitzman et al., 2012; Howe et al., 2018; Aitken et al., 2016; Stockdale et al., 2019), often due to fears that the entity might violate consent terms by sharing data with unauthorized third parties (Kacsmar et al., 2022; Middleton et al., 2020). This distrust is intensified when participants are unfamiliar with the recipient entity (Weitzman et al., 2012; Howe et al., 2018), especially if it is a governmental or for-profit organization (Aitken et al., 2016; Stockdale et al., 2019; Garrison et al., 2016; Hendricks-Sturrup and Lu, 2023; Khanra, 2023; American Medical Association, 2022). Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling (Seh et al., 2020). People are also reluctant to interact with entities with any history of data breaches, doubting their ability to safeguard sensitive information (Seh et al., 2020).
@@
Mental models on PETs.Mental models on PETsMental models on PETs. PETs aim to address user privacy concerns. While various PETs are available for health data donation environments (Jordan et al., 2022), prior work indicates that participants with greater online privacy literacy tend to have more trusting beliefs in PETs (Harborth and Pape, 2020). Studies focusing on sensitive personal data, including health data, reveal similar findings. For instance, Dechand et al. (Dechand et al., 2019) found that most participants do not trust encryption because they do not understand how it works: e.g., many believe that service providers can bypass encryption to gain unauthorized data access. Furthermore, Lerner et al. (Lerner et al., 2022) noted that some participants doubt the existence of true privacy.
Non-experts may misunderstand privacy technologies (Wu and Zappala, 2018; Oates et al., 2018; Dechand et al., 2019), e.g., viewing encryption merely as restrictive access control (Wu and Zappala, 2018) or confusing it with data encoding (Abu-Salma et al., 2017). Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections (Tang et al., 2021; Velykoivanenko et al., 2022; Abu-Salma et al., 2018; Gerber et al., 2019; Pham and Castro, 2019; Vaniea et al., 2014). As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality. We investigate the alignment between stated privacy guarantees and people’s expectations for how their data will be protected as part of RQ1.
@@
Explaining PETs.Explaining PETsExplaining PETs. To address these misunderstandings and evaluate the impact of PETs on downstream factors such as privacy concerns (or expectations) and willingness to share data, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts (Cranor et al., 2006). Prior work focuses heavily on E2EE and DP (Abu-Salma et al., 2018; Bai, 2019; Cummings et al., 2021; De Luca et al., 2016; Akgul et al., 2021; Dechand et al., 2019; Abu-Salma et al., 2017; Wu and Zappala, 2018; Xiong et al., 2020; Nanayakkara et al., 2023). These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge. Methods found effective to explain PETs include visualizations (Stransky et al., 2021), mental models (Stewart and Lacey, 2012; Lin et al., 2012), nutrition labels (Kelley et al., 2010), metaphors (Karegar et al., 2022), short statements (Papacharissi and Fernback, 2005), and privacy games (Tupsamudre et al., 2018; Karl and Tao, 2023). In this work we use short statements to briefly explain the relevant guarantees in order to evaluate the impact of these guarantees on participants’ privacy expectations and willingness to donate medical data.
@@
Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing. Prior work on the effect of PETs and their accompanying PGs typically focuses on the effect of a combination of guarantees offered by a specific PET (e.g., DP (Valdez and Ziefle, 2019; Cummings et al., 2021) or E2EE (Dechand et al., 2019)) on privacy concerns . A more limited subset of studies evaluate the impact on willingness to share data (Valdez and Ziefle, 2019).
@@
Instead of focusing on a specific PET and it’s combined accompanying PGs we seek to examine and compare four PGs that are shared across many PETs: data expiration, data anonymization, use restriction, access control, and auditing (see § 1). In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity. Furthermore, we focus on the effect of two different auditing processes – expert auditing and self auditing – on enhancing privacy expectations and willingness to share data. Other studies have examined public understanding of data retention (Murillo et al., 2018) and general data anonymization (Haddow et al., 2011), and one study examines willingness to donate medical data if it is allowed to be used for secondary purposes (Juga et al., 2021). To our knowledge, no prior work focuses on comparing the effects of the four core PGs we study nor the effects of auditing PETs on user privacy expectations or willingness to donate data.
3. Methodology
@@
As mentioned in § 1, we address research questions RQ1 and RQ2 through a user survey. In this section, we discuss the ethical considerations of our study, outline the survey design (§ 3.2), and describe the cognitive interview process used to ensure participants’ understanding of the presented scenarios (§ 3.3). We also describe our participant selection strategy (§ 3.4) and the limitations of our methodology (§ 3.6).
3.1. Ethics
Before we began our research, our ethics review board approved all of the study methods, including the cognitive interviews and the survey. A consent form detailing the data storage, handling, and survey compensation is presented at the beginning of the survey. Participants can opt out of disclosing any demographic data without influencing their compensation in the survey.

### v3

@@
Abstract
While voluntary donation of private health information enables valuable research, privacy concerns often deter potential donors. Privacy Enhancing Technologies (PETs) aim to address these concerns, yet their effectiveness in encouraging data sharing remains unclear. This study conducts a vignette survey ($N=494$) with participants recruited from Prolific to examine the willingness of US-based people to donate medical data for developing new treatments. It investigates four general guarantees offered across PETs: data expiration, anonymization, purpose restriction, and access control and two mechanisms for verifying these guarantees: self-auditing and expert auditing. This study also controls for the impact of confounds, including demographics and two types of data collectors: for-profit and non-profit institutions.
@@ Abstract
I Introduction
@@
The altruistic use of personal health data holds immense potential for societal benefit. For example, the public health sector can leverage individuals’ medical histories and health data to analyze epidemiological trends, enhance disease surveillance, improve risk prediction models, diagnose rare and emerging diseases, and accelerate the development of novel treatments [1, 2, 3, 4, 5, 6].
Beyond direct medical applications, health data sharing offers economic benefits. A unified data-sharing infrastructure could eliminate redundant clinical trials and research efforts, thereby reducing healthcare costs [7]. According to the European Commission, innovations that enhance health data accessibility could generate savings of billions of euros within the European healthcare sector alone [8]. These demands are further intensified by recent advances in artificial intelligence, particularly machine learning models, whose effectiveness heavily depends on the quality and quantity of data they are trained on.
Recognizing these converging medical, scientific, and economic imperatives, the National Academy of Medicine (NAM) has advocated for more open and collaborative approaches to health data sharing [9], emphasizing both the ethical obligation and practical necessity of broader data access frameworks.
However, despite these compelling benefits and institutional support for expanded health data sharing, privacy concerns remain a significant barrier to widespread data donation. Personal health data is routinely collected by healthcare providers [3, 5], mobile applications, and wearable devices [4], often flowing to third parties through data-sharing agreements [6] with limited transparency to the individuals who generated it. Studies consistently show that privacy concerns are a primary deterrent to individuals contributing their sensitive health information [10, 11, 12, 4, 5], particularly when hospitals permit data sharing with third parties without explicit patient consent [13].
In an effort to foster data collection and analysis while protecting the privacy of data contributors, there has been an increasing focus on deploying privacy-enhancing technologies (PETs) for data storage [14, 15], data processing [16, 17, 18, 19, 20, 21, 22], and machine learning (ML) training [23, 24]. Some systems explicitly consider data donation scenarios: Waldo [25] enables privacy-preserving queries over medical time-series data for remote patient monitoring; CoVault [24] supports expressive queries on sensitive data at scale under a strong threat model, and considers a national-scale epidemic analytics scenario; Anonify [22] provides decentralized dual-level anonymity for medical data donation; and Mycelium [26] supports differentially-private queries over large distributed graphs (e.g., disease-spread data) across millions of user devices.
However, despite these technological advances, implementing PETs alone does not guarantee increased data availability. Privacy protections must be expressed to potential donors in comprehensible terms that address their fundamental concerns, or these sophisticated mechanisms remain ineffective at encouraging data sharing.
@@
In order to evaluate whether the guarantees offered by PETs ultimately impact people’s willingness to donate their data, at least three steps are necessary. First, we must be able to effectively explain guarantees (either general guarantees or those of a particular PET) to the people they aim to protect: end-users. Prior work has focused on doing so for particular PETs, mainly differential privacy (DP) and end-to-end encryption (E2EE) [27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Second, using these explanations, we must evaluate whether the guarantees impact people’s privacy concerns (or expectations), and in turn, their willingness to share data. Prior work has done such evaluations using self-report studies focused on individual PETs, again primarily DP [27, 37, 36] or E2EE [32, 38, 28]. Third, while rare, we must validate those results in the field to confirm that the effects observed replicate when studying actual versus intended behavior (e.g., [39, 40]).
In this work, we focus on addressing the first two steps outlined above: effectively explaining privacy guarantees and evaluating their impact on data-sharing intentions. We follow real-world medical consent frameworks [41, 42] and privacy regulations [43, 44], and individually investigate four fundamental guarantees across different technologies: anonymization, access control, data expiration, and purpose restriction (PG(1)–PG(4)), along with two verification processes: expert and self auditing (AG(1)–AG(2)). This approach helps identify which specific protections most influence privacy expectations and data-sharing decisions.
@@
We examine the role of these PET guarantees and their audits in shaping privacy expectations and influencing data-sharing intentions within a specific context used in prior work (see e.g., [27, 36]): medical data donation. We ask:
@@ 1. RQ(1):
    How well do people understand and expect what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?
@@ 2. RQ(2):
To address these questions, we conduct a vignette survey ($n=494$) following the best practice methodology in prior work [27, 28, 35]. Each survey respondent is presented with a hypothetical opportunity to donate their health data to help develop a treatment for a specific chronic disease, along with how their data will be protected (PG(1)–PG(4) enforced (or not) by AG(1)–AG(2)). We control for confounding factors identified in prior work, including data-collection entity [6, 45] and socio-demographics [46, 47, 6].
@@
Furthermore, while the technical community has suggested external audits as a mechanism to increase trust in PET implementation, our initial inquiry suggests that more work is needed to explain the purpose and effectiveness of such audits to end-users. In fact, the effect of audit statements on people’s willingness to donate is limited to a specific scenario involving for-profit entities and auditing to check that purpose restriction (PG(4)) is correctly implemented.
We argue that it is critical for non-profit entities to rigorously implement data privacy measures , as any future data leak could lead to a significant loss of trust. Prior research [40] underscores that users place greater value on maintaining their expected privacy than on gaining additional privacy they did not initially anticipate. In contrast, we highlight the risk of for-profit entities engaging in “privacy washing” [48], where statements about PETs are used to artificially raise privacy expectations to encourage data collection. At the same time, our findings reveal that respondents are perceptive of the limitations in general PET guarantees, particularly concerning protections against data breaches. This underscores the need for future research to explore how to effectively communicate the potential of stronger, emerging PET guarantees and address the skepticism of end-users.
II Related Work
In this section, we examine prior work on people’s willingness to donate personal data, privacy concerns, the role of PETs, and educational or explanatory strategies for PETs.
@@ II Related Work
Data donation.Data donationData donation. The analysis of personal health data is crucial for medical research, particularly in diagnosing emerging or rare diseases and developing new treatments. Indeed, the COVID-19 pandemic has further intensified the demand for personal health data [49]. Despite the importance of this data, the sensitive nature of health data poses significant obstacles to its collection [50]. Prior work indicates that individuals are generally willing to donate their data for altruistic purposes [51, 2, 52], though there is a notable reluctance when it comes to their health data specifically [53, 54]. This reluctance diminishes when the donor or their close family members are directly affected by the disease under study [55, 3], suggesting that non-privacy-related factors influence donation decisions.
@@
Privacy concerns.Privacy concernsPrivacy concerns. Concerns about privacy and the misuse of donated data are prominent among potential donors [54]. These concerns include the risk of being identified, discriminated against, and having personal sensitive data misused or leaked [56, 2, 5, 57]. A common source of these concerns is distrust of the receiving entity [56, 58, 2, 59], often due to fears that the entity might share data with unauthorized third parties [60, 53]. This distrust is intensified when participants are unfamiliar with the recipient entity [56, 58], or if it is a governmental or for-profit organization [2, 59, 61, 45, 62, 63]. Beyond this intentional data misuse, there are concerns about data leakage caused by hackers or unintentional mishandling [64]. People are reluctant to interact with entities having any history of data breaches, doubting their ability to safeguard sensitive information [64].
@@
Mental models on PETs.Mental models on PETsMental models on PETs. PETs aim to address user privacy concerns by providing technical safeguards for sensitive data. While various PETs are available for health data donation contexts [65], their effectiveness depends largely on users’ trust and understanding. Prior work indicates that participants with greater online privacy literacy tend to have more trusting attitudes in PETs [66], while those with limited knowledge often remain skeptical.
Studies have also found that non-experts systematically misunderstand privacy technologies [34, 67, 32]. For instance, many incorrectly conceptualize encryption as merely a form of access control [34] or confuse it with simple data encoding [33]. Additionally, non-expert end users may struggle to understand the consequences of inadequate privacy protections [68, 69, 28, 70, 71, 72]. Lerner et al. [73] identified an even more fundamental barrier: some participants express inherent skepticism about the existence of “true” privacy, illustrating how deep-seated misconceptions can fundamentally undermine the reassurance PETs are designed to provide.
As a result, even when privacy protections are present, people may have risk expectations that are misaligned with reality. We investigate the alignment between stated privacy guarantees and people’s expectations for how their data will be protected as part of RQ1.
@@
Explaining PETs.Explaining PETsExplaining PETs. To address these misunderstandings and evaluate the impact of PETs on downstream factors such as privacy concerns (or expectations) and willingness to share data, technologists and researchers have worked to explain PETs to the public to address their unfamiliarity with privacy concepts [74]. Prior work focuses heavily on E2EE and DP [28, 30, 27, 31, 29, 32, 33, 34, 35, 36]. These efforts remain ongoing, as effectively and scalable setting privacy expectations remains a challenge. Methods found effective to explain PETs include visualizations [75], mental models [76, 77], nutrition labels [78], metaphors [79], short statements [80], and privacy games [81, 82]. However, there is a gap in the literature regarding techniques to explain auditing guarantees, despite the importance of auditing in verifying compliance with privacy promises.
@@
Privacy guarantees (PGs) and auditing.Privacy guarantees (PGs) and auditingPrivacy guarantees (PGs) and auditing. Prior research on PETs and their accompanying PGs has primarily focused on evaluating specific technologies (e.g., DP [83, 27] or E2EE [32]) and their combined impact on privacy concerns. A smaller subset of studies has assessed how these technologies affect willingness to share data [83]. Additional research has examined public understanding of specific guarantees such as data retention [84], data anonymization [85], and secondary use permissions [86].
@@
Our approach differs by systematically comparing four core PGs implemented across many PETs: data expiration, data anonymization, use restriction, and access control (see § I). We also separately investigate auditing mechanisms, which serve as verification procedures rather than direct guarantees. In particular, we examine whether offering a given PG influences people’s willingness to donate health data to the recipient entity. Furthermore, we focus on the effect of two different auditing processes – expert auditing and self auditing – on enhancing privacy expectations and willingness to share data.
To our knowledge, this is the first study to systematically compare these four core PGs and evaluate how different auditing mechanisms affect user privacy expectations and willingness to donate health data.
III Methodology
@@
As mentioned in § I, we address research questions RQ1 and RQ2 through a user survey. In this section, we discuss the ethical considerations of our study (§ III-A), outline the explored statements (§ III-B) and survey design (§ III-C), and describe the cognitive interview and pilot study process used to refine and improve participants’ comprehension of the presented scenarios (§ III-D). We also describe our participant selection strategy (§ III-E), analysis procedure (§ III-F), and the limitations of our methodology (§ III-G).
III-A Ethics
Our study was conducted under approval from our university’s Ethics Review Board (ERB). The approval covered both the cognitive interviews and the survey. We implemented several measures to ensure ethical treatment of participants:
All survey data was collected anonymously through Qualtrics [87], with IP address collection disabled. Participants were presented with a consent form at the beginning of the survey that detailed the study purpose, data handling procedures, and compensation details. Participants could opt out of answering any demographic questions without affecting their compensation.
For cognitive interviews conducted via Zoom [88], participants were asked to avoid using their real names when joining the meetings to maintain anonymity. Only anonymized transcripts were retained for analysis after removing any personal identifiable information.
All research participants were recruited anonymously through Prolific [89], providing an additional layer of separation between researchers and participants’ identities. The collected data was accessible only to the research team, and all data will be deleted after the completion of the research project in accordance with our ERB protocol.